John Billiard Supply sells all types of billiard equipment, and is considering
manufacturing their own brand of pool cues. John Farris, the production manager,
determines that the materials and labor cost for each cue is $30 and the fixed cost
that must be covered is $2700 per week. If the selling price for each cue is $50.
Then
a. How many pool cues must be sold per week to breakeven point?
b. What would be the new selling price per unit for a break-even of 300 units?
Answer #1

Solution:
Given in the question
Selling price = 50
Fixed cost = 2700
Labour cost = 30
Lets assume that X quantity is for breakeven point
So for breakeven Cost = Sell
2700 + 30*X = 50*X
20*X = 2700
X = 135
So we need 135 pool cues must be sold per week to breakeven point.

Solution(b)
if Quantity is 300 units than assume that new selling price is X.
300*X = 2700 + 30*300
300*X = 2700+9000 = 11700
X = 39
So new selling price should be $39